# Seal2Real: Prompt Prior Learning on Diffusion Model for Unsupervised Document Seal Data Generation and Realisation

### v2

@@ Abstract
Seal-related tasks in document processing—such as seal segmentation, authenticity verification, seal removal, and text recognition under seals—hold substantial commercial importance. However, progress in these areas has been hindered by the scarcity of labeled document seal datasets, which are essential for supervised learning. To address this limitation, we propose Seal2Real, a novel generative framework designed to synthesize large-scale labeled document seal data. As part of this work, we also present Seal-DB, a comprehensive dataset containing 20,000 labeled images to support seal-related research. Seal2Real introduces a prompt prior learning architecture built upon a pre-trained Stable Diffusion model, effectively transferring its generative capability to the unsupervised domain of seal image synthesis. By producing highly realistic synthetic seal images, Seal2Real significantly enhances the performance of downstream seal-related tasks on real-world data. Experimental evaluations on the Seal-DB dataset demonstrate the effectiveness and practical value of the proposed framework.
@@ 1 Introduction
Seals are commonly present in Chinese documents such as contracts, agreements, statements, and notices. With the ongoing digitization of documents, an increasing number are scanned into images for digital feature extraction. However, the presence of seals significantly impacts various downstream document recognition and processing tasks 1, 2, 3, 4. For instance, seals can interfere with accurate text recognition. Consequently, numerous seal-related tasks have emerged in document processing 5, 6, 7, including seal segmentation, authenticity discrimination, and text recognition beneath seals—tasks of considerable commercial importance.
@@
To address these limitations, synthesizing large-scale seal image datasets has become essential. However, a critical requirement is that the synthetic data must be sufficiently realistic to support downstream applications by closely mimicking the appearance of real seals in scanned documents. In summary, a crucial upstream task is the generation of seal images. Given a document image $I_{d}$ without a seal, the objective is to generate a forged version $I_{f}$ with a synthetic seal, where the position, size, and text content of the seal in $I_{f}$ can be manually customized, as illustrated in Fig. 1. This approach enables the automatic generation of corresponding labels.
@@
Traditional approaches to seal synthesis 8, 9, 10, 11, 12, 13, 14, 15 typically rely on geometric correction, template matching, and color distribution 5. However, these methods often suffer from limited robustness and require complex parameter tuning, making it difficult to generate visually convincing forged seals. Specifically, they often fail to capture the subtle textures, ink bleed variations, paper interactions, and degradation artifacts common in real-world scanned or photographed seals. Their effectiveness in achieving high-fidelity seal synthesis remains constrained. With the advancement of deep learning in computer vision 16, modern generative models—such as GANs 17, 18, 19, 20 and Diffusion Models 21, 22, 23, 24, 25, 26, 27—offer more powerful alternatives for high-fidelity image generation.
@@
In this paper, we propose Seal2Real, an unsupervised method for seal image generation and realization, leveraging a pretrained Stable Diffusion Model 28, along with a large-scale dataset, Seal-DB, comprising both real and synthetic seal images. Our primary contributions are as follows:
@@
(2) We design a seal forger network that enhances the realism of synthetic seals, guided by the learned priors, allowing them to closely mimic real ones. This not only augments the dataset but also empowers downstream tasks—such as seal authenticity discrimination—to effectively identify forged seals.
@@ 2.2 Prior Learning Stage
In the first stage, we aim to learn two distinct prompts: a real prompt, referring to real seal images, and a forgery (or fake) prompt, referring to all synthetic seal images. These prompts are learned using the pretrained Stable Diffusion Model and its simple loss function 29, applied to our proposed Seal-DB dataset. As shown in Fig. 5, Seal-DB comprises two parts: synthetic seal and real seal document images. The synthetic seal images are generated using our traditional seal synthesis method described in Sec. 1.
The initial real and forgery prompts serve as coarse textual descriptions of real and fake seal images, as illustrated in Fig. 4(a). Specifically, we initialize the real prompt as “A photo of document with real seal,” with its corresponding embedding $T_{r}\in\mathbb{R}^{N\times 768}$, and the forgery prompt as “A photo of document with fake seal,” with embedding $T_{f}\in\mathbb{R}^{N\times 768}$ 30, where $N$ denotes the length of each prompt.
Given a real seal image $I_{r}$, we input $I_{r}$ and the real prompt $T_{r}$ into the pretrained Stable Diffusion Model $SD_{\theta}(I_{r},t_{r},T_{r})$. The image $I_{r}$ is first encoded into a latent representation $z_{r}$, to which random Gaussian noise $\epsilon_{r}$ is added, resulting in the noised latent $z_{r,t_{r}}$. The model $SD_{\theta}(I_{r},t_{r},T_{r})$ then estimates the added noise $\hat{\epsilon}_{r}$, where $\theta$ represents the parameters of the U-Net 31 in the Stable Diffusion Model 28. Similarly, for a synthetic seal image $I_{s}$, we input $I_{s}$ and the forgery prompt $T_{f}$ into $SD_{\theta}(I_{s},t_{f},T_{f})$.
It is important to note that the noise terms $\epsilon_{r},\epsilon_{f}$ and time steps $t_{r}$, $t_{f}$ are randomly sampled at each iteration to approximate the law of large numbers. Based on the text-conditioned noise estimation loss 32 in latent space, we apply the simple loss $\mathcal{L}_{prompts}$ to learn the real and forgery prompt pair while keeping the U-Net parameters $\theta$ frozen.
@@ 2.3 Forger Network Learning Stage
After learning the real and forgery prompts in the first stage, we proceed to train a forger network using a prompt-prior loss. As a baseline, we adopt a simple U-Net architecture 31 for our forger $f_{\phi}$ to enhance the realism of synthetic seal images and to validate the effectiveness of our framework.
@@
While Eq. 3 plays a critical role in guiding the forger network to generate images aligned with the learned prompts—thereby enhancing the realism of the forgery—it may not sufficiently preserve the texture and structural details of the input image, potentially resulting in a completely different visual output. Conversely, Eq. 4 is essential for maintaining the content and structure of the original synthetic seal image. However, relying solely on this loss may fail to achieve a convincingly realistic appearance in the forged result.
@@ 2.4 Training with Frozen Encoder
In our setup, the CLIP text encoder is kept frozen, meaning its parameters remain unchanged throughout training. This is essential for maintaining consistent text representations, which enables effective control over the generation of real and synthetic seal images based on textual prompts. During training, real prompts are paired with actual seal images, while fake prompts are associated with synthetic seal images. This dichotomous pairing ensures that the fine-tuned Stable Diffusion model learns to distinguish effectively between the two underlying distributions. As a result, by leveraging distinct textual prompts, our model is capable of generating outputs that accurately reflect the intended nature—real or synthetic—of the input prompt.
@@ 3.2 Analysis of Experimental Results
To further demonstrate the necessity of our method, Fig. 6 presents results from several state-of-the-art text-to-image generation models, including PixArt-$\alpha$ 33, SD-XL, SDXL-turbo 34, Wuerstechen 35, and Kandinsky-3 36, along with an ablation study using a GAN-based baseline. These comparisons clearly illustrate the superiority and robustness of employing a pretrained diffusion model over training a GAN from scratch.
@@
Image
ID
Synthetic Prompt Type
The Forger
Network
GAN
@@
Fig. 8 illustrates the visual differences in synthetic images before and after applying the forger network. In addition, we compare our Seal2Real framework with several state-of-the-art text-conditioned image editing methods, as such methods might appear suitable for this task. Specifically, we evaluate InstructP2P 37, EditDDPM 38, and MGIE 39, using the editing instruction: “Make this seal more like the real one.”
@@
For each task, we adopt established baseline methods and modify only the training datasets to assess the impact of our synthesized data. Specifically, for seal segmentation, we use the U-Net architecture 31; for authenticity classification, we employ a CLIP-based approach 40; and for OCR in documents containing seals, we utilize SwinTextSpotter 41. All evaluations are conducted using the same quantity of training data, comparing the original dataset with our generated dataset to highlight the improvements enabled by our method.

### v3

@@ Abstract
Seal-related tasks in document processing-such as seal segmentation, authenticity verification, seal removal, and text recognition under seals-hold substantial commercial importance. However, progress in these areas has been hindered by the scarcity of labeled document seal datasets, which are essential for supervised learning. To address this limitation, we propose Seal2Real, a novel generative framework designed to synthesize large-scale labeled document seal data. As part of this work, we also present Seal-DB, a comprehensive dataset containing 20,000 labeled images to support seal-related research. Seal2Real introduces a prompt prior learning architecture built upon a pre-trained Stable Diffusion model, effectively transferring its generative capability to the unsupervised domain of seal image synthesis. By producing highly realistic synthetic seal images, Seal2Real significantly enhances the performance of downstream seal-related tasks on real-world data. Experimental evaluations on the Seal-DB dataset demonstrate the effectiveness and practical value of the proposed framework. The dataset is available at https://github.com/liuyifan6613/DocBank-Document-Enhancement-Dataset.
@@ 1 Introduction
Seals are commonly present in Chinese documents such as contracts, agreements, statements, and notices. With the ongoing digitization of documents, an increasing number are scanned into images for digital feature extraction. However, the presence of seals significantly impacts various downstream document recognition and processing tasks 1, 2, 3, 4. For instance, seals can interfere with accurate text recognition. Consequently, numerous seal-related tasks have emerged in document processing 5, 6, 7, including seal segmentation, authenticity discrimination, and text recognition beneath seals-tasks of considerable commercial importance.
@@
To address these limitations, synthesizing large-scale seal image datasets has become essential. Similar synthetic data generation and AI-augmented pipelines have demonstrated significant potential in addressing data scarcity within related fields, such as 3D content creation and immersive AR interaction 8, 9, 10. However, a critical requirement is that the synthetic data must be sufficiently realistic to support downstream applications by closely mimicking the appearance of real seals in scanned documents. In summary, a crucial upstream task is the generation of seal images. Given a document image $I_{d}$ without a seal, the objective is to generate a forged version $I_{f}$ with a synthetic seal, where the position, size, and text content of the seal in $I_{f}$ can be manually customized, as illustrated in Fig. 1. This approach enables the automatic generation of corresponding labels.
@@
Traditional approaches to seal synthesis 11, 12, 13, 14, 15, 16, 17, 18 typically rely on geometric correction, template matching, and color distribution 5. However, these methods often suffer from limited robustness and require complex parameter tuning, making it difficult to generate visually convincing forged seals. Specifically, they often fail to capture the subtle textures, ink bleed variations, paper interactions, and degradation artifacts common in real-world scanned or photographed seals. Their effectiveness in achieving high-fidelity seal synthesis remains constrained. With the advancement of deep learning in computer vision 19, modern generative models-such as GANs 20, 21, 22, 23 and Diffusion Models 24, 25, 26, 27, 28, 29, 30-offer more powerful alternatives for high-fidelity image generation. These architectures have recently pushed the boundaries of complex visual reasoning in domains like spatio-temporal autonomous driving 31 and vision-language navigation 32.
@@
In this paper, we propose Seal2Real, an unsupervised method for seal image generation and realization, leveraging a pretrained Stable Diffusion Model 33, along with a large-scale dataset, Seal-DB, comprising both real and synthetic seal images. Our primary contributions are as follows:
@@
(2) We design a seal forger network that enhances the realism of synthetic seals, guided by the learned priors, allowing them to closely mimic real ones. This not only augments the dataset but also empowers downstream tasks-such as seal authenticity discrimination-to effectively identify forged seals.
@@ 2.2 Prior Learning Stage
In the first stage, we aim to learn two distinct prompts: a real prompt, referring to real seal images, and a forgery (or fake) prompt, referring to all synthetic seal images. These prompts are learned using the pretrained Stable Diffusion Model and its simple loss function 34, applied to our proposed Seal-DB dataset. As shown in Fig. 5, Seal-DB comprises two parts: synthetic seal and real seal document images. The synthetic seal images are generated using our traditional seal synthesis method described in Sec. 1.
The initial real and forgery prompts serve as coarse textual descriptions of real and fake seal images, as illustrated in Fig. 4(a). Specifically, we initialize the real prompt as “A photo of document with real seal,” with its corresponding embedding $T_{r}\in\mathbb{R}^{N\times 768}$, and the forgery prompt as “A photo of document with fake seal,” with embedding $T_{f}\in\mathbb{R}^{N\times 768}$ 35, where $N$ denotes the length of each prompt.
Given a real seal image $I_{r}$, we input $I_{r}$ and the real prompt $T_{r}$ into the pretrained Stable Diffusion Model $SD_{\theta}(I_{r},t_{r},T_{r})$. The image $I_{r}$ is first encoded into a latent representation $z_{r}$, to which random Gaussian noise $\epsilon_{r}$ is added, resulting in the noised latent $z_{r,t_{r}}$. The model $SD_{\theta}(I_{r},t_{r},T_{r})$ then estimates the added noise $\hat{\epsilon}_{r}$, where $\theta$ represents the parameters of the U-Net 36 in the Stable Diffusion Model 33. Similarly, for a synthetic seal image $I_{s}$, we input $I_{s}$ and the forgery prompt $T_{f}$ into $SD_{\theta}(I_{s},t_{f},T_{f})$.
It is important to note that the noise terms $\epsilon_{r},\epsilon_{f}$ and time steps $t_{r}$, $t_{f}$ are randomly sampled at each iteration to approximate the law of large numbers. Based on the text-conditioned noise estimation loss 37 in latent space, we apply the simple loss $\mathcal{L}_{prompts}$ to learn the real and forgery prompt pair while keeping the U-Net parameters $\theta$ frozen.
@@ 2.3 Forger Network Learning Stage
After learning the real and forgery prompts in the first stage, we proceed to train a forger network using a prompt-prior loss. As a baseline, we adopt a simple U-Net architecture 36 for our forger $f_{\phi}$ to enhance the realism of synthetic seal images and to validate the effectiveness of our framework.
@@
While Eq. 3 plays a critical role in guiding the forger network to generate images aligned with the learned prompts-thereby enhancing the realism of the forgery-it may not sufficiently preserve the texture and structural details of the input image, potentially resulting in a completely different visual output. Conversely, Eq. 4 is essential for maintaining the content and structure of the original synthetic seal image. However, relying solely on this loss may fail to achieve a convincingly realistic appearance in the forged result.
@@ 2.4 Training with Frozen Encoder
In our setup, the CLIP text encoder is kept frozen, meaning its parameters remain unchanged throughout training. This is essential for maintaining consistent text representations, which enables effective control over the generation of real and synthetic seal images based on textual prompts. During training, real prompts are paired with actual seal images, while fake prompts are associated with synthetic seal images. This dichotomous pairing ensures that the fine-tuned Stable Diffusion model learns to distinguish effectively between the two underlying distributions. As a result, by leveraging distinct textual prompts, our model is capable of generating outputs that accurately reflect the intended nature-real or synthetic-of the input prompt.
@@ 3.2 Analysis of Experimental Results
To further demonstrate the necessity of our method, Fig. 6 presents results from several state-of-the-art text-to-image generation models, including PixArt-$\alpha$ 38, SD-XL, SDXL-turbo 39, Wuerstechen 40, and Kandinsky-3 41, along with an ablation study using a GAN-based baseline. These comparisons clearly illustrate the superiority and robustness of employing a pretrained diffusion model over training a GAN from scratch.
@@
Image ID Synthetic Prompt Type The Forger Network GAN
@@
Fig. 8 illustrates the visual differences in synthetic images before and after applying the forger network. In addition, we compare our Seal2Real framework with several state-of-the-art text-conditioned image editing methods, as such methods might appear suitable for this task. Specifically, we evaluate InstructP2P 42, EditDDPM 43, and MGIE 44, using the editing instruction: “Make this seal more like the real one.”
@@
For each task, we adopt established baseline methods and modify only the training datasets to assess the impact of our synthesized data. Specifically, for seal segmentation, we use the U-Net architecture 36; for authenticity classification, we employ a CLIP-based approach 45; and for OCR in documents containing seals, we utilize SwinTextSpotter 46. All evaluations are conducted using the same quantity of training data, comparing the original dataset with our generated dataset to highlight the improvements enabled by our method.